General Meeting Summary 10/18/18

From Cuawiki
Jump to: navigation, search

MAGNET (Bogdan)

  • not yet closed as still need to attach upstream coil
  • expect to be fully assembled this week


HV BASES (Julie)

  • Assembly and testing is ongoing


FRAME AND PROTOTYPE TESTS (Carlos)

  • Working on optimizing construction and minimizing material between crystals - prototype tests ongoing
  • Current idea for assembly is to assemble from the back and use rectangular plastic clips/frames at the front to hold assembly together
  • Tests being performed with prototype include mechanical assembly procedure, thermal tests, and perhaps magnetic field tests
  • Discussion about cabling plans
  • based on Frascati experience with PDME, a cable chicane may not be necessary - a light tight channel may be sufficient, but it is important that it also has air flow
  • multiwire cables are generally rigid and might be difficult in assembly
  • overall idea is to get cables our of box, then at a distance of ~1m have a transition box to multiwire cables and then these would go wherever needed
  • Bogdan can send a reference for HV connectors used in Hall A calorimeter/coordinate detector
  • IPN-O will do a few tests with multiwire cables to test structure - note that box size should be as small as possible
  • Discussion about infrastructure outside box - further discussion in November with engineers/designers at JLab


READOUT (Alexandre)

  • Component are mostly available modulo 1-2 additional VXS crates
  • Test setup in SHMS for cables and background tests
  • Plan is to start on beam test setup during the next extended down period
  • availability of 3x3 crystal prototype: hardware currently disassembled, could be reassembled in a few days
  • monitoring/curing system hardware also available


CRYSTALS AND PROTOTYPE (Vladimir, Sasha)

  • currently installing light monitoring system
  • Plan is to move detector to Hall D next week Tuesday and power up to check all channels - platform should be available in November
  • First five crystals from Crytur are being tested - quality seems very good. Next six to be shipped before end of the month. Discussion about future procurement ongoing


NPS COLLABORATION MEETING

  • Defer until next meeting when perhaps dates of Hall A/C winter meetings are known


NEXT MEETING: THURSDAY 8 NOVEMBER at 8:00AM (EST)